The following information is for Punta Company for July: Factory overhead costs were applied to jobs at the predetermined rate of $42.50 per labor hour. Job S incurred 6,175 labor hours; Job T used 4,275 labor hours. Job S was shipped to customers during July. Job T was still in process at the end of July. The overapplied or underapplied overhead to the Cost of Goods Sold account was closed at the end of July. Factory utilities, factory depreciation, and factory insurance incurred are summarized as follows: Utilities $ 14,250 Depreciation 45,000 Insurance 18,000 Total $ 77,250 f. Direct materials and indirect materials used are as follows: Job S Job T Total Material A $ 28,500 $ 71,250 $ 99,750 Material B 12,000 35,000 47,000 Subtotal $ 40,500 $ 106,250 $ 146,750 Indirect materials 211,000 Total $ 357,750 g. Direct labor incurred for the two jobs and indirect labor are as follows: Job S $ 55,500 Job T 45,000 Indirect labor 133,000 Total $ 233,500 Required: 1. Calculate the

total manufacturing cost for Job S and Job T for July. (Round your intermediate calculations and final answers to 2 decimal places.) 2. Calculate the amount of overapplied or underapplied overhead and state whether the Cost of Goods Sold account will be increased or decreased by the adjustment.
Answer #1

Solution 1:

Job S Job T
Direct material 40500.00 106250.00
Direct labor 55000.00 45000.00
overhead Applied (Hours* $42.50) 262437.50 181687.50
Total manufacturing cost 357937.50 332937.50

Solution 2:

Actual Overhead = $77250 + $211000 + $133000 = $421,250

Applied overhead = 262437.50 + 181687.50 = $444,125

Overapplied overhead = Applied overhead - Actual Overhead = 444,125 - 421,250 = $22,875 Overapplied

Cost of Goods Sold account will be decreased by $22,875.